# Josephson effect in a Fibonacci quasicrystal

### v1

@@ II.3 Superconductivity and Josephson junction
Next, we estimate the effective superconducting coherence length in our model system. We consider the commonly used expression for the ballistic superconducting coherence length $\xi_{0}=\hbar{v_{\mathrm{F}}}/(\pi\Delta_{0})$, with Fermi velocity ${v_{\mathrm{F}}}$ computed from the normal-state dispersion in the leads (using a single-band model at half-filling). We obtain $\xi_{0}\approx 11a_{0}$. This value is slightly smaller but still the same order as obtained through fitting the exponential decay of the ABS in the main text (see Sec. III.1), where we find $\xi_{0}\approx 17a_{0}$. We use these estimates as a representative length scale to quantify the normal junction length $L$, and for simplicity assume that $L$ is the same as the effective junction length [138], such that a long (short) Josephson junction corresponds to $L\gg\xi_{0}$ ($L\ll\xi_{0}$). Thus, the short junction limit $L\ll\xi_{0}$ essentially corresponds to a quantum dot ($L\sim a_{0}$) in our system. We note that the interesting physics comes from the unique quasiperiodic spatial dependence, while such a quantum dot has no spatial extent. Still, for full transparency we consider junctions from single sites to thousands of sites, covering both the short and long junction regimes.
@@ III Spectrum and current-phase relation
We start by studying the DC Josephson effect through a Fibonacci chain by solving Eq. (1) and computing the resulting subgap ABS spectrum ($E<|\Delta|$) and the current-phase relation. We find that the broken mirror symmetry in the quasiperiodic lattice breaks the well-known ABS degeneracy points in the spectrum, while the current-phase relation always shows a traditional $2\pi$ period with either sinusoidal or sawtooth shape. We show how quasiperiodicity can enhance the current at lower interface transmission, and demonstrate that the low-energy ABS, which carry most of the current, obtain quasiperiodic oscillations. We also comment on the influence of microscopic even-odd effects, which strongly influence the degeneracy points in the ABS spectrum and subsequently the value of the current.
III.1 ABS spectrum and current-phase relation
In order to highlight the influence of quasiperiodicity on the Josephson effect, we in this subsection study how the low-energy spectrum and Josephson current depend on the superconducting phase difference $\Delta\theta$ for different hopping ratios $\rho$ and interface hoppings ${t_{\mathrm{int}}}$, keeping other parameters fixed. We study the subgap ABS which drive most of the current, and demonstrate how the current can be enhanced by quasiperiodicity. Finally, we study the probability density of the ABS, which obtain quasiperiodic oscillations on top of the otherwise resonant level.
@@ III.1 ABS spectrum and current-phase relation
Figure 3 shows the low-energy spectrum (a) and Josephson current-phase relation (b) in a Josephson junction where the non-superconducting region is the Fibonacci approximant $C_{9}$ for different $\rho$ (line colors) at ${t_{\mathrm{int}}}=0.7{t_{\mathrm{B}}}$, ${\rho_{\mathrm{S}}}=1$ and $\mu=0$. We note that all energy levels in Fig. 3(a) are subgap ABS and that the number of such states is directly related to e.g. superconducting gap versus level spacing (or analogously the superconducting coherence length versus junction length) [151]. Notably, the crystalline junction ($\rho=1$) shows the usual ABS degeneracy at $\Delta\theta=0$ and $2\pi$, while the quasiperiodic junction ($\rho\neq 1$) shows no such degeneracy. Generally, the degeneracy can break whenever there is an asymmetry between the two leads [151]. Indeed, we find that the broken degeneracy is caused by the quasiperiodic hopping structure $C_{9}={t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}{\bm{{t_{\mathrm{A}}}}}\ldots{\bm{{t_{\mathrm{B}}}}}{t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}$ which breaks mirror symmetry whenever $\rho\neq 1$, thus causing a different local hopping neighborhood at each superconducting lead. We verify that the degeneracy is also broken in other asymmetric approximants e.g. $C_{10}={t_{\mathrm{A}}}\ldots{t_{\mathrm{B}}}$, while it is conserved in symmetric approximants e.g. $C_{3}={t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}$ regardless of $\rho$ (or number of repetitions $N$). More generally, we also find conserved degeneracy when the characteristic function Eq. (2) is terminated at a point which conserves mirror symmetry (e.g. when the number of bonds are $F_{n}-2$). These results illustrate the important concept of how different Fibonacci chains can lead to both qualitatively and quantitatively different behavior, a reoccurring theme throughout our work.
@@
Next, we focus on the Josephson current in Fig. 3(b), which in this case is enhanced by quasiperiodicity for $\rho<1$. We can explain this in terms of how the Andreev reflection is modified by the quasiperiodic hopping asymmetry $\rho\neq 1$, together with the superconducting hopping ratio ${\rho_{\mathrm{S}}}$ and interface hopping ${t_{\mathrm{int}}}$, since they effectively model the transmission together. To start, it is well-known that a crystalline junction shows maximal current at ideal junction transmission (i.e. at perfect Andreev reflection and thus zero normal reflection), where the current-phase relation turns into a linear sawtooth profile with a discontinuity at $\Delta\theta=\pi$ [140]. At this phase difference, the lowest-energy ABS becomes an exact degenerate zero-energy state with perfect resonance (i.e. wave-function matching) [141], also associated with a Jackiw-Rebbi zero mode [152]. But this perfect scenario is in a sense fine-tuned since any deviation from ideal transmission lifts the zero-energy degeneracy [141]. It is reasonable to expect that realistic materials show a reduced transmission due to interface imperfections or other effects. Specifically, Fig. 3 corresponds to a system with noticeably lower transmission for the crystalline junction $\rho=1$ (as modeled by ${t_{\mathrm{int}}}=0.7{t_{\mathrm{B}}}$). For such reduced junction transmission, there is an increased occurrence of normal reflection instead of Andreev reflection, and the resonance is not perfect. As a result, the ABS is shifted to finite energy, obtaining a smaller and non-linear phase dispersion, thus yielding a reduced current [153] with the sinusoidal shape seen for $\rho=1$ in Fig. 3(b). The quasiperiodic junction, on the other hand, instead approaches large current and sawtooth profile for smaller values of $\rho$. We note that this can also occur for larger $\rho$ depending on the other model parameters (which we show in Sec. IV). Thus, the quasiperiodic hopping asymmetry modifies the ABS spectrum and condition for an ABS at zero energy (i.e. due to perfect Andreev reflection), the latter we denote ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ and which we fully quantify for different quasiperiodic junctions in Sec. IV.
@@
III.2 Microscopic even-odd effect
Before diving deeper into the influence of quasiperiodicity on the current, we first have to comment on the strong influence of microscopic even-odd effects on the spectrum, specifically in a junction with even or odd number of energy levels. The scenario also illustrates what happens in a system where there is already a zero-energy state before the onset of superconductivity. Depending on the actual physical realization of the system, this might either be an idealized or a crucial effect.
In Sec. III.1 we showed how the ABS in the crystalline junction ($\rho=1$) are degenerate at $\Delta\theta=0$ and $2\pi$, and how the degeneracy is lifted for all $\rho\neq 1$ when the quasiperiodic junction breaks the mirror symmetry across the junction. The Fibonacci approximant $C_{9}$ studied in that case has an even number of sites $F_{9}+1=56$, while the number of sites instead becomes odd when repeating the approximant an even number of times, e.g. $1112\times F_{9}+1=$ 111, or in every third approximant e.g. $C_{8}$ with $F_{8}+1=35$ sites. The normal-state spectrum with such an odd number of sites (and thus energy levels) must have an energy level at exactly zero energy due to chiral symmetry, and the ABS spectrum becomes qualitatively modified with the ABS degeneracy points shifting to $\Delta\theta=\pi$ instead [151]. We verify that this occurs also in quasiperiodic junctions as illustrated in Fig. 5, showing the energy-phase spectrum (a) and current-phase relation (b) for a junction consisting of a repeated $C_{9}$ approximant (i.e. otherwise equivalent to Fig. 3). In Fig. 5(a) quasiperiodicity breaks the degeneracy at $\Delta\theta=\pi$ at all higher energies due to the broken mirror symmetry, but the degeneracy at zero energy is maintained leading to a fully linear phase dispersion with only small variations in slope with $\rho$ at low energy. Consequently, since this zero-energy state carries most of the current, $\rho$ therefore has negligible influence on the current, causing a similar sawtooth profile in all cases shown in Fig. 5(b).
Next, we find that at the critical phase difference $\Delta\theta=\Delta{\theta_{\mathrm{c}}}$, the current is completely carried by the lowest-energy state, with the current contribution from all other states are either exactly zero or pair-wise cancelling. Away from this phase, the reduction in current comes primarily from the variation in the energy-phase dispersion from a linear slope at $\Delta\theta=\pi$ to zero slope at $\Delta\theta=0$ or $2\pi$, and secondarily from a destructive contribution of higher-energy states. This behavior is qualitatively different from the junction with even number of sites described in Sec. III.1, where the energy-phase dispersion changes also its functional form away from $\Delta{\theta_{\mathrm{c}}}$. Importantly, the zero-energy state in Fig. 5(a) is robust against variations in the hopping parameters. We find that this is related to a preserved degeneracy where e.g. the probability density of the zero-energy state is exactly the same in both leads (not shown), in contrast to the higher-energy states which break the degeneracy.
@@ IV Critical current
Having established the influence of quasiperiodicity on the ABS spectrum and current-phase relation in Sec. III.1, we next look to its influence on observable features like the critical current. In particular, we quantify how the critical current ${I_{\mathrm{c}}}$ depend on the hopping parameters in different approximants $C_{n}$ with and without repetition, and in more general quasiperiodic junctions of length $L$ as modeled by the characteristic function in Eq. (2). We further determine the functional form of the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ for the emergence of a zero-energy ABS, which we use to analyze several emergent features including the behavior of the critical current as function of junction length.
@@ IV.1 Zero-energy ABS and critical current
We start by studying the dependence of the spectrum and critical current on ${t_{\mathrm{int}}}$ and $\rho$. In particular, Fig. 7 shows the energy eigenvalue of the ABS closest to zero energy ${E_{\mathrm{min}}}$ (a) and the critical current ${I_{\mathrm{c}}}$ (b) as functions of ${t_{\mathrm{int}}}$ for different $\rho$ (line colors), in a junction described by the Fibonacci approximant $C_{9}$. Here, it is important to note that the regime ${t_{\mathrm{int}}}>{t_{\mathrm{B}}}$ may be as reasonable as ${t_{\mathrm{int}}}<{t_{\mathrm{B}}}$ since there are two hopping terms in the Fibonacci chain (${t_{\mathrm{A}}}$ and ${t_{\mathrm{B}}}$) and another in the superconducting leads (${t_{\mathrm{S}}}$), which might be quite dissimilar. In particular, we verify that the most important results in the following discussion are valid when e.g. ${t_{\mathrm{B}}}<{t_{\mathrm{int}}}\leq{t_{\mathrm{S}}},{t_{\mathrm{A}}}$. In Fig. 7(a) ${E_{\mathrm{min}}}$ approaches zero at a specific ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ due to perfect Andreev reflection [141], correlating with the maximal critical current in Fig. 7(b). We find that the current is fully carried by the lowest-energy state at these extrema due to zero contribution from all other states, and that the current-phase relation becomes a sawtooth profile (not shown). In contrast, away from the extrema in ${E_{\mathrm{min}}}$ and ${I_{\mathrm{c}}}$, the energy-phase slope reduces and the current-phase relation becomes sinusoidal, where other states also start contributing destructively. Interestingly, Fig. 7 illustrates that for lower transmission ${t_{\mathrm{int}}}<{t_{\mathrm{B}}}$, quasiperiodicity can significantly enhance the current for most $\rho<1$, without any parameter fine-tuning.
@@
To understand the behavior in ${E_{\mathrm{min}}}$ and ${I_{\mathrm{c}}}$, we simultaneously vary $\rho$, ${\rho_{\mathrm{S}}}$ and ${t_{\mathrm{int}}}$ and find that ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})/{t_{\mathrm{B}}}=\rho\sqrt{{\rho_{\mathrm{S}}}}$ for the $C_{9}$ approximant, see Appendix A for calculation details. This is consistent with the positions of the extrema (e.g. ${t_{\mathrm{int}}^{*}}/{t_{\mathrm{B}}}\approx 1$ for $\rho={\rho_{\mathrm{S}}}=1$) up to a small deviation which we attribute to corrections in powers of $|\Delta|/{t_{\mathrm{B}}}$, i.e. vanishing for $|\Delta|\ll{t_{\mathrm{B}}},{t_{\mathrm{A}}},{t_{\mathrm{S}}}$ [141]. We proceed to quantify the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ in Table 1 for different Fibonacci approximants up to $C_{17}$ where the junction length $L=a_{0}F_{n}$ is more than two (three) orders of magnitude larger than the coherence length $\xi_{0}$ (atomic scale $a_{0}$). We find that in junctions with even number of sites, ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ is described by one of three functional forms $\rho\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{{\rho_{\mathrm{S}}}}$, thus always scaling as $\sqrt{{\rho_{\mathrm{S}}}}$, but where the scaling in $\rho$ varies. Thus, we find qualitatively different behavior from in crystalline junctions, where it is known that a normal region with a single hopping ${t_{\mathrm{A}}}$ scales as ${t_{\mathrm{int}}^{*}}({t_{\mathrm{A}}},{t_{\mathrm{S}}})=\sqrt{{t_{\mathrm{A}}}{t_{\mathrm{S}}}}$ in the limit $|\Delta|\ll{t_{\mathrm{A}}},{t_{\mathrm{S}}}$ [141]. Furthermore, in the junctions with odd number of sites (i.e. every third approximant), there is already a zero-energy ABS in the normal state [151], and we here let ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ denote for which $\rho$ and ${\rho_{\mathrm{S}}}$ the state remains at zero energy. Thus, we find a zero-energy ABS $\forall{t_{\mathrm{int}}},{\rho_{\mathrm{S}}}$ for $\rho=1$ (and $\forall\rho,{\rho_{\mathrm{S}}}$ at ${t_{\mathrm{int}}}\to 0$) as described in Sec. III.2.
@@
Figure 8 shows the critical current as a function of ${t_{\mathrm{int}}}$ through the $C_{10}$ junction with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$ (a) and through the $C_{12}$ junction with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{{\rho_{\mathrm{S}}}}$ (b). These cases show smaller or no spread in ${I_{\mathrm{c}}}$ with $\rho$, respectively, as compared to $C_{9}$ with linear scaling $\rho\sqrt{{\rho_{\mathrm{S}}}}$ in Fig. 7(b). We also note that there is still a variation in the peak values for different $\rho$ in Fig. 8(b), which we find to be caused by how $\rho$ changes the slope of the energy-phase dispersion. Furthermore, we point out that quasiperiodicity can still influence the spread of the peaks in the $C_{12}$, since it still varies with ${t_{\mathrm{B}}}$, but here we keep ${t_{\mathrm{B}}}$ fixed, as the unit of energy. Thus, allowing the hopping ${t_{\mathrm{B}}}$ to vary can lead to ${t_{\mathrm{int}}^{*}}({t_{\mathrm{A}}},{t_{\mathrm{B}}},{t_{\mathrm{S}}})$ being proportional to e.g. ${t_{\mathrm{B}}}$ or $\sqrt{{t_{\mathrm{B}}}}$.
@@
Next, we briefly summarize how the condition for zero-energy ABS behaves in more generalized quasiperiodic junctions modeled by the characteristic function in Eq. (2), for different lengths $L\in[2,200]a_{0}$ (see Appendix A for calculation details). We find that junctions with odd number of sites follow the same behavior as either Fig. 9(a) or (b), i.e. with a robust zero-energy mode $\forall\rho$, or only when $\rho=1$, respectively. In junctions with even number of sites, we find that the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ always scales as $\propto\sqrt{{\rho_{\mathrm{S}}}}$, while the scaling in $\rho$ varies between different rational exponents, e.g. $\rho^{3/2}$, $\rho$, $\sqrt{\rho}$, $1/\sqrt{\rho}$ and $1/\rho$. Interestingly, we find that as $L$ increases, the system varies between the conditions in a quasiperiodic manner similar to how the hopping structure evolves with $L$. This is consistent also with the approximants $C_{n}$, where the change in the functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ is cyclic in $n$, and therefore quasiperiodic in $L$ following the Fibonacci numbers, since $L=a_{0}F_{n}$.

### v2

@@ II.3 Superconductivity and Josephson junction
Next, we estimate the effective superconducting coherence length in our model system. We consider the commonly used expression for the ballistic superconducting coherence length $\xi_{0}=\hbar{v_{\mathrm{F}}}/(\pi\Delta_{0})$, with Fermi velocity ${v_{\mathrm{F}}}$ computed from the normal-state dispersion in the leads (using a single-band model at half-filling). We obtain $\xi_{0}\approx 11a_{0}$. This value is slightly smaller but still the same order as obtained through fitting the exponential decay of the ABS in the main text (see Sec. III), where we find $\xi_{0}\approx 17a_{0}$. We use these estimates as a representative length scale to quantify the normal junction length $L$, and for simplicity assume that $L$ is the same as the effective junction length [143], such that a long (short) Josephson junction corresponds to $L\gg\xi_{0}$ ($L\ll\xi_{0}$). Thus, the short junction limit $L\ll\xi_{0}$ essentially corresponds to a quantum dot ($L\sim a_{0}$) in our system. We note that the interesting physics comes from the unique quasiperiodic spatial dependence, while such a quantum dot has no spatial extent. Still, for full transparency we consider junctions from single sites to thousands of sites, covering both the short and long junction regimes.
@@ III Spectrum and current-phase relation
In order to highlight the influence of quasiperiodicity on the DC Josephson effect, we in this section study how the ABS spectrum ($E<|\Delta|$) and Josephson current depend on the superconducting phase difference $\Delta\theta$ for different hopping ratios $\rho$ and interface hoppings ${t_{\mathrm{int}}}$, keeping other parameters fixed.
@@
Figure 3 shows the low-energy spectrum (a) and Josephson current-phase relation (b) in a Josephson junction where the non-superconducting region is the Fibonacci approximant $C_{9}$ for different $\rho$ (line colors) at ${t_{\mathrm{int}}}=0.7{t_{\mathrm{B}}}$, ${\rho_{\mathrm{S}}}=1$ and $\mu=0$. We note that all energy levels in Fig. 3(a) are subgap ABS and that the number of such states is directly related to e.g. superconducting gap versus level spacing (or analogously the superconducting coherence length versus junction length) [158]. Notably, the crystalline junction ($\rho=1$) shows the usual ABS degeneracy at $\Delta\theta=0$ and $2\pi$, while the quasiperiodic junction ($\rho\neq 1$) shows no such degeneracy. Generally, the degeneracy can break whenever there is an asymmetry between left- and right-moving quasiparticle states, for instance related to an asymmetry between the two leads [158]. Here, we find that the broken degeneracy in the quasiperiodic junction is related to the $C_{9}={t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}{\bm{{t_{\mathrm{A}}}}}\ldots{\bm{{t_{\mathrm{B}}}}}{t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}$ approximant having different local hopping neighborhood at each superconducting lead whenever $\rho\neq 1$. We verify that the degeneracy is also broken in other asymmetric approximants e.g. $C_{10}={t_{\mathrm{A}}}\ldots{t_{\mathrm{B}}}$, while it is conserved in symmetric approximants e.g. $C_{3}={t_{\mathrm{A}}}{t_{\mathrm{B}}}{t_{\mathrm{A}}}$ regardless of $\rho$ (or number of repetitions $N$). More generally, we also find conserved degeneracy when the characteristic function Eq. (2) is terminated at a point which conserves the symmetry between leads (e.g. when the number of bonds are $F_{n}-2$). These results illustrate the important concept of how different Fibonacci chains can lead to both qualitatively and quantitatively different behavior, a reoccurring theme throughout our work.
@@
Next, we focus on the Josephson current in Fig. 3(b), which in this case is enhanced by quasiperiodicity for all $\rho<1$ compared to the crystalline case ($\rho=1$). We can explain this in terms of how the Andreev reflection is modified by the quasiperiodic hopping asymmetry $\rho\neq 1$, together with the superconducting hopping ratio ${\rho_{\mathrm{S}}}$ and interface hopping ${t_{\mathrm{int}}}$, since they effectively model the transmission together. To start, it is well-known that a crystalline junction shows maximal current at ideal junction transmission (i.e. at perfect Andreev reflection and thus zero normal reflection), where the current-phase relation turns into a linear sawtooth profile with a discontinuity at $\Delta\theta=\pi$ [145]. At this phase difference, the lowest-energy ABS becomes an exact degenerate zero-energy state with perfect resonance (i.e. wave-function matching) [146], also associated with a Jackiw-Rebbi zero mode [159]. But this perfect scenario is in a sense fine-tuned since any deviation from ideal transmission lifts the zero-energy degeneracy [146]. It is reasonable to expect that realistic materials show a reduced transmission due to interface imperfections or other effects. Specifically, Fig. 3 corresponds to a system with noticeably lower transmission for the crystalline junction $\rho=1$ (as modeled by ${t_{\mathrm{int}}}=0.7{t_{\mathrm{B}}}$). For such reduced junction transmission, there is an increased occurrence of normal reflection instead of Andreev reflection, and the resonance is not perfect. As a result, the ABS is shifted to finite energy, obtaining a smaller and non-linear phase dispersion, thus yielding a reduced current [160] with the sinusoidal shape seen for $\rho=1$ in Fig. 3(b). The quasiperiodic junction, on the other hand, instead approaches large current and sawtooth profile for smaller values of $\rho$. We note that this can also occur for larger $\rho$ depending on the other model parameters (which we show in Sec. IV). Thus, the quasiperiodic hopping asymmetry modifies the ABS spectrum and condition for an ABS at zero energy (i.e. due to perfect Andreev reflection), the latter we denote ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ and which we fully quantify for different quasiperiodic junctions in Sec. IV.
@@
Finally, we comment that the ABS degeneracy and current-phase relation studied above can be significantly altered in crystalline junctions by microscopic effects, e.g. with a zero-energy state existing already in the normal state due to symmetry [158], which can result in a sawtooth current-phase profile for all interface transparencies. In Appendix A we investigate these effects in the presence of quasiperiodicity, and find that such a robust sawtooth profile is either maintained or broken depending on the quasiperiodic approximant structure.
@@ IV Critical current
Having established the influence of quasiperiodicity on the ABS spectrum and current-phase relation in Sec. III, we next look to its influence on observable features like the critical current. In particular, we here quantify how the critical current ${I_{\mathrm{c}}}$ defined in Eq. (7) depends on the hopping parameters in different approximants $C_{n}$ with and without repetition, and in more general quasiperiodic junctions of length $L$ as modeled by the characteristic function in Eq. (2). We further determine the functional form of the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ for the emergence of a zero-energy ABS, which we use to analyze several emergent features. Finally, we compute the critical current as function of junction length and find that it exhibits quasiperiodic oscillations superposed on the usual decay found in crystalline junctions. While some of the parameter dependencies considered in this section may not be tunable in solids (although they might be in certain atomically engineered chains or synthetic quasiperiodic metamaterials [161, 162, 163, 164, 165]), our theoretical model calculations serve to establish the fundamental influence of the quasiperiodic hopping modulation on the critical current.
@@ IV.1 Zero-energy ABS and critical current
We start by studying the dependence of the spectrum and critical current on ${t_{\mathrm{int}}}$ and $\rho$. In particular, Fig. 5 shows the energy eigenvalue of the ABS closest to zero energy ${E_{\mathrm{min}}}$ (a) and the critical current ${I_{\mathrm{c}}}$ (b) as functions of ${t_{\mathrm{int}}}$ for different $\rho$ (line colors), in a junction described by the Fibonacci approximant $C_{9}$. Here, it is important to note that the regime ${t_{\mathrm{int}}}>{t_{\mathrm{B}}}$ may be as reasonable as ${t_{\mathrm{int}}}<{t_{\mathrm{B}}}$ since there are two hopping terms in the Fibonacci chain (${t_{\mathrm{A}}}$ and ${t_{\mathrm{B}}}$) and another in the superconducting leads (${t_{\mathrm{S}}}$), which might be quite dissimilar. In particular, we verify that the most important results in the following discussion are valid when e.g. ${t_{\mathrm{B}}}<{t_{\mathrm{int}}}\leq{t_{\mathrm{S}}},{t_{\mathrm{A}}}$. In Fig. 5(a) ${E_{\mathrm{min}}}$ approaches zero at a specific ${t_{\mathrm{int}}}={t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ due to perfect Andreev reflection [146], correlating with the maximal critical current in Fig. 5(b). We find that the current is fully carried by the lowest-energy state at these extrema due to zero contribution from all other states, and that the current-phase relation becomes a sawtooth profile (not shown). In contrast, away from the extrema in ${E_{\mathrm{min}}}$ and ${I_{\mathrm{c}}}$, the energy-phase slope reduces and the current-phase relation becomes sinusoidal, where other states also start contributing destructively. Interestingly, Fig. 5 illustrates that for ${t_{\mathrm{int}}}<{t_{\mathrm{B}}}$ (${t_{\mathrm{int}}}>{t_{\mathrm{B}}}$), quasiperiodicity can significantly enhance the current for most $\rho<1$ ($\rho>1$) as compared to the crystalline junction with $\rho=1$, i.e. without any parameter fine-tuning.
@@
To understand the behavior in ${E_{\mathrm{min}}}$ and ${I_{\mathrm{c}}}$, we simultaneously vary $\rho$, ${\rho_{\mathrm{S}}}$ and ${t_{\mathrm{int}}}$ and find that ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})/{t_{\mathrm{B}}}=\rho\sqrt{{\rho_{\mathrm{S}}}}$ for the $C_{9}$ approximant, see Appendix B for calculation details. This is consistent with the positions of the extrema (e.g. ${t_{\mathrm{int}}^{*}}/{t_{\mathrm{B}}}\approx 1$ for $\rho={\rho_{\mathrm{S}}}=1$) up to a small deviation which we attribute to corrections in powers of $|\Delta|/{t_{\mathrm{B}}}$, i.e. vanishing for $|\Delta|\ll{t_{\mathrm{B}}},{t_{\mathrm{A}}},{t_{\mathrm{S}}}$ [146]. We proceed to quantify the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ in Table 1 for different Fibonacci approximants up to $C_{17}$ where the junction length $L=a_{0}F_{n}$ is more than two (three) orders of magnitude larger than the coherence length $\xi_{0}$ (atomic scale $a_{0}$). We find that in junctions with even number of sites, ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ is described by one of three functional forms $\rho\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$, $\sqrt{{\rho_{\mathrm{S}}}}$, thus always scaling as $\propto\sqrt{{\rho_{\mathrm{S}}}}$. Thus, we find qualitatively different behavior from in crystalline junctions, where it is known that a normal region with a single hopping ${t_{\mathrm{A}}}$ scales as ${t_{\mathrm{int}}^{*}}({t_{\mathrm{A}}},{t_{\mathrm{S}}})=\sqrt{{t_{\mathrm{A}}}{t_{\mathrm{S}}}}$ in the limit $|\Delta|\ll{t_{\mathrm{A}}},{t_{\mathrm{S}}}$ [146]. Furthermore, in the junctions with odd number of sites (i.e. every third approximant), there is already a zero-energy ABS in the normal state [158], and we here let ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ denote for which $\rho$ and ${\rho_{\mathrm{S}}}$ the state remains at zero energy. Thus, we find a zero-energy ABS $\forall{t_{\mathrm{int}}},{\rho_{\mathrm{S}}}$ for $\rho=1$ (and $\forall\rho,{\rho_{\mathrm{S}}}$ at ${t_{\mathrm{int}}}\to 0$) as described in Appendix A.
@@
Figure 6 shows the critical current as a function of ${t_{\mathrm{int}}}$ through the $C_{10}$ junction with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{\rho}\sqrt{{\rho_{\mathrm{S}}}}$ (a) and through the $C_{12}$ junction with ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})=\sqrt{{\rho_{\mathrm{S}}}}$ (b). These cases show smaller or no spread in ${I_{\mathrm{c}}}$ with $\rho$, respectively, as compared to $C_{9}$ with linear scaling $\rho\sqrt{{\rho_{\mathrm{S}}}}$ in Fig. 5(b). We also note that there is still a variation in the peak values for different $\rho$ in Fig. 6(b), which we find to be caused by how $\rho$ changes the slope of the energy-phase dispersion. Furthermore, we point out that quasiperiodicity can still influence the spread of the peaks in the $C_{12}$, since it still varies with ${t_{\mathrm{B}}}$, but here we keep ${t_{\mathrm{B}}}$ fixed, as the unit of energy. Thus, allowing the hopping ${t_{\mathrm{B}}}$ to vary in the model calculations can lead to ${t_{\mathrm{int}}^{*}}({t_{\mathrm{A}}},{t_{\mathrm{B}}},{t_{\mathrm{S}}})$ being proportional to e.g. ${t_{\mathrm{B}}}$ or $\sqrt{{t_{\mathrm{B}}}}$.
@@
Next, we briefly summarize how the condition for zero-energy ABS behaves in more generalized quasiperiodic junctions modeled by the characteristic function in Eq. (2), for different lengths $L\in[2,200]a_{0}$ (see Appendix B for calculation details). We find that junctions with odd number of sites follow the same behavior as either Fig. 7(a) or (b), i.e. with a robust zero-energy mode $\forall\rho$, or only when $\rho=1$, respectively. In junctions with even number of sites, we find that the condition ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ always scales as $\propto\sqrt{{\rho_{\mathrm{S}}}}$, while the scaling in $\rho$ varies between different rational exponents, e.g. $\rho^{3/2}$, $\rho$, $\sqrt{\rho}$, $1/\sqrt{\rho}$ and $1/\rho$. Interestingly, we find that as $L$ increases, the system varies between the conditions in a quasiperiodic manner similar to how the hopping structure evolves with $L$. This is consistent also with the approximants $C_{n}$, where the change in the functional form of ${t_{\mathrm{int}}^{*}}(\rho,{\rho_{\mathrm{S}}})$ is cyclic in $n$, and therefore quasiperiodic in $L$ following the Fibonacci numbers, since $L=a_{0}F_{n}$.